# PanoNormal: Monocular Indoor 360° Surface Normal Estimation

### v1

@@ Abstract
The presence of spherical distortion on the Equirectangular image is an acknowledged challenge in dense regression computer vision tasks, such as surface normal estimation. Recent advances in convolutional neural networks (CNNs) strive to mitigate spherical distortion but often fall short in capturing holistic structures effectively, primarily due to their fixed receptive field. On the other hand, vision transformers (ViTs) excel in establishing long-range dependencies through a global self-attention mechanism, yet they encounter limitations in preserving local details. We introduce PanoNormal, a monocular surface normal estimation architecture designed for 360° images, which combines the strengths of CNNs and ViTs. Specifically, we employ a multi-level global self-attention scheme with the consideration of the spherical feature distribution, enhancing the comprehensive understanding of the scene. Our experimental results demonstrate that our approach achieves state-of-the-art performance across multiple popular 360° monocular datasets. The code and models will be released.
@@ 1 Introduction
Figure 1: Normal estimation results on real-world instances and previously unseen data from SUN360 [23]. Our approach, when compared to (b) PanoFormer [18] and (c) HyperSphere [14], stands out for its ability to provide precise surface normal details, even in challenging scenarios.
@@
Estimating surface normal is a crucial task in computer vision as it provides essential geometric information about the structure and orientation of surfaces within a scene. By offering valuable insights into the underlying 3D geometry of objects, accurate surface normal estimation contributes significantly to advancing the capabilities of a wide range of applications, including object recognition, autonomous driving, and robotics. Despite significant progress in normal estimation for conventional perspective images  [4, 16], the task of 360° normal estimation remains less explored. When directly applying the methods proposed for perspective images to 360° images, the results are unsatisfactory. This is because the spatial warping inherent in 360° images is not taken into account in these techniques and so we instead, in this paper, develop specialized techniques tailored to their spherical representation.
To solve the distortions in panoramas in other similar tasks such as 360° depth estimation, some works have proposed projection-fusion architectures such as UniFuse [13] to leverage complementary information from the equirectangular projection (ERP) image and cubemap projection (CMP) patches to refine prediction outcomes in the spherical domain. However, this strategy needs to bridge the domain gap between different projections, and the additional cross-projection fusion module introduces extra computational overhead. Distortion-aware deep learning frameworks are also proposed to mitigate such issues [8, 16], embedding spherical geometry explicitly or implicitly into CNNs, or devising specialized loss functions tailored for panoramic images. More recently, researchers have adopted self-attention mechanisms to recapture a holistic understanding of the entire scene within the 360° domain [15, 18]. While CNNs demonstrate prowess in localized feature extraction, they exhibit limitations in preserving the holistic structure of the 360° scene. Conversely, transformers guarantee the modelling of long-range dependencies but show shortcomings in capturing fine-grained details. Furthermore, conventional transformer-based methods cannot reveal the inherent spherical properties of the input signal. These limitations of CNNs and transformers led us to investigate hybrid architectures that seamlessly integrate the strengths of both for 360° images, fostering a refined and comprehensive representation of the scene across diverse scales for 360° surface normal estimation.
@@
In this paper, we present PanoNormal, a framework for monocular surface normal estimation in 360° indoor environments. PanoNormal strategically integrates the strengths of CNNs for extracting low-level features and enhancing locality, along with the advantages of transformers in capturing and associating long-range dependencies. We introduce a deep learning module aimed at extracting low-level features from raw 360° images, departing from the conventional approach of direct tokenization inputs. To enhance the capabilities of self-attention in exploring the structure of the surface normal map, we propose a multi-scale transformer decoder, enriching the final representation by integrating information across diverse scales. In Fig. 1, previous state-of-the-art methods deliver sub-optimal results when applied to real-world scenes from SUN360. In contrast, our approach predicts sharper and more accurate details of local areas and finely-delineated object boundaries, significantly improving the holistic geometry understanding of the scene. We perform comprehensive experiments on public datasets to assess the effectiveness of our approach compared to state-of-the-art models. Our contributions are summarized as follows:
@@
  • Extensive experiments have conducted on widely recognized benchmarks, namely 3D60, Stanford2D3D, Matterport3D, SunCG, and Structured3D. The results illustrate that PanoNormal consistently surpasses state-of-the-art approaches, and contributes to further studies in this domain by providing the first comprehensive evaluation across these public benchmarks.
@@ 2 Related Work
Most existing methods [9, 5, 7] for surface normal estimation are designed for perspective images, which poses challenges, such as distortions, deformations, and domain shifts, that make conventional methods unsuitable for 360° images. Moreover, panoramic images introduce additional complexities to surface normal estimation, including variable illumination and intricate indoor layouts. Recent works can be broadly categorized into direct and indirect methods. Direct methods [16, 14, 11] introduce specific loss functions (e.g., spherical loss, hyper-sphere loss, double-quaternion loss) to comprehend the geometric structure on the unit sphere with various CNN architectures. [8] explicitly embed invariance against spherical distortions into CNNs, adjusting the sampling locations of convolutional filters to counter distortions and align the filters around the sphere effectively. Indirect methods [2, 27] draw inspiration from other vision tasks, e.g., 360° depth estimation, and adapt them for surface normal estimation. Work such as [15, 13, 21, 18] propose methodologies that project equirectangular images into corresponding CMP or tangent plane projection (TP) sub-images for predictions or directly learn internal correlations among these projections, aiming to capture details and holistic information concurrently with distortion awareness. However, employing simple CNNs with various spherical loss functions or adapting surface normal estimation methods from depth alone fails to yield satisfactory results. In addition, there are no surface normal estimation methods that provide a standardized evaluation across diverse public benchmarks, similar to the established practices in depth estimation tasks.
@@ 2.2 Vision Transformer
Vision transformers (ViTs) are a class of models that use self-attention mechanisms to process images as sequences of patches [10]. ViTs [27, 22, 12, 29] have achieved remarkable results on conventional images captured by pinhole cameras, which have a limited field of view and which have, at most, mild distortions. In contrast, 360° images offer a complete, immersive view of the surrounding scene but require adaptations and modifications of the original ViTs to handle challenges such as significant image distortions, object deformation, and domain shifts. [18] and [15] proposed to divide patches on the spherical tangent domain into tangent tokens to reduce the negative effect of panoramic distortions, and introduced an attention module that considers both spatial and angular relations among tokens. [25] presented a panoramic ViT for saliency detection in 360° videos, which leverages a multi-scale feature fusion module and a temporal attention module to capture the spatial and temporal saliency cues. [26] proposed a method for panoramic semantic segmentation task, which is equipped with deformable patch embedding and deformable MLP modules for handling object deformations and image distortions, and also enhances the mutual prototypical adaptation strategy for unsupervised domain adaptive panoramic segmentation. These works demonstrate the potential and effectiveness of ViTs for 360° vision tasks and motivated us to investigate suitable ViTs for the surface normal dense prediction task, aiming to seamlessly integrate the strengths of the previously mentioned methods. Our objective is to yield robust evaluation results across benchmarks, contributing to further studies in this domain.
@@ 3 Our Method
To address the challenge of estimating surface normals using a single panoramic image for indoor scenes, we propose a novel architecture, PanoNormal. PanoNormal leverages the capabilities of convolutional layers for extracting meaningful embeddings from panoramic images. To address the challenge of spherical distortion in 360° images while maintaining global dependencies, we crafted a U-shaped distortion-aware transformer architecture, drawing inspiration from the approach proposed by [18]. More specifically, we adopt a multi-level structure in our transformer decoder, capturing spatial relationships, processing fine-grained details, and integrating high-level contextual information across different levels. Fig. 2 illustrates the architecture of PanoNormal.
@@ 3.1.1 Feature Embedding
In a conventional ViT-based network, feature embedding undergoes processing through only a single convolutional layer before ViT blocks. It has been demonstrated that it works well for some 360° understanding tasks, such as depth estimation [18]. Nonetheless, although both depth estimation and surface normal tasks demand high-quality contextual information, surface normal tasks necessitate deeper features to effectively capture more useful cues, such as the orientation of objects. In contrast to the simplicity suitable for depth estimation, the intricate nature of surface normal requirements poses a challenge for the feature embedding achieved through a single convolutional layer, leading to sub-optimal results (quantitative results are shown in Tab. 1). To address this issue, we introduce a series of convolutional layers aimed at extracting more effective features. This approach enhances our model’s ability to gain a more comprehensive understanding of the entire scene, improving performance in surface normal estimation. The feature embedding block consists of three $3\times 3$ convolutional layers, each followed by a batch normalization function and a rectified linear unit (ReLU) activation function.
@@ 3.1.2 Distortion-aware Transformer Encoder
The inherent non-uniform spatial warping of visual features, caused by spherical distortion in the 360° image, is alleviated by applying PanoFormer blocks, as introduced by [18]. To address issues such as distortions, misalignment, and inaccuracies arising from direct pixel sampling on the ERP image, a tangent projection (TP) division strategy is employed. It samples related surrounding pixels of the central pixel (tangent point) on each dense divided plane in the tangent domain, obtaining a series of tokens in patches with decent position representation that can be further transformed to the ERP domain. Subsequently, a conventional multi-head vision transformer block is applied, with the replacement of the feed-forward network (FFN) to the locally-enhanced feed-forward network (LeFF) [24, 22] for enhancing local feature interaction. The introduced encoder not only computes attention scores between central and correlated tokens but also integrates a trainable token flow. This token flow serves as a bias to adjust the spatial distribution of tokens, supplying the network with additional positional information for learning global dependencies. An example is shown in Fig. 2(a) and the representation of the self-attention is as follows:
@@
where the sampling strategy is adopted on the feature representations $f$, and is denoted as $\hat{s}$; $m,q,k$ represent the self-attention head, each token, and its neighboring tokens in a tangent patch, respectively. $W_{m}$ and $W^{\prime}_{m}$ denote the learnable weights of each head, $A_{mqk}$ is the attention weights for each token, and $\Delta s_{mqk}$ signifies the learned bias through of each token’s query.
@@ 3.1.3 Multi-level Transformer Decoder
Our proposed ViT decoder, designed with spherical distortion awareness and a hierarchical structure to predict normal maps of different scales, is illustrated in Fig. 2(b). The proposed multi-level decoder comprises four independent blocks. Each block processes the concatenation of the upsampled encoded representations $\hat{f}_{i}$, featuring twice the spatial resolution size and halved the number of channels from the preceding level, and the directly propagated features $\hat{s}_{i}$ from the encoders through the skip-links. The surface normal vectors in various scales are generated as outputs using $3\times 3$ convolutions, followed by a hyperbolic tangent (tanh) activation function to constrain them to the range $[-1,1]$. This process can be described as:
@@ 3.2 Loss Function
The presented network produces surface normal maps at various scales to capture a more comprehensive global geometry structure of the scene. In the training process, the predicted maps apply bilinear upsampling to align with the input size, followed by the adoption of the following losses:
3.2.1 MSE Loss:
$L_{m}$, is the mean squared error between the ground truth and the predicted normals of each pixel, defined as:
@@ 3.2.1 MSE Loss:
3.2.2 Quaternion Loss:
$L_{q}$ [14], measures the angular difference between predicted and ground truth normal maps on a pixel-wise basis:
@@ 3.2.2 Quaternion Loss:
where $M$ indicates the number of pixels of the input, and $j$ indexes the current pixel.
3.2.3 Perceptual Loss:
$L_{p}$, is employed on the finest scales to enhance the generation of finer details:
$L_{p}=l_{feat}^{\phi,k}(\hat{\textbf{N}},\textbf{N})=\sum_{j=1}^{M}\frac{1}{C_{k}M}\|\phi_{k}(\hat{\textbf{N}}_{j})-\phi_{k}(\textbf{N}_{j})\|_{2}^{2}$ (5)
where $\phi$ is the VGG-16 network [19] that pretrained on the ImageNet dataset [17], and $C_{k}$ indicates $C$ dimensional features for the $k$-th layer of the network $\phi$.
3.2.4 Smooth Loss:
$L_{s}$, quantifies the gradient $G$ in the $x$ and $y$ directions in the ground truth and the predicted surface normal map at all scales:
@@ 3.2.4 Smooth Loss:
By default, we set $\lambda_{m}=1.0$, $\lambda_{q}=10.0$, $\lambda_{p}=0.05$, and $\lambda_{s}=0.5$ as the weights for different terms. In our experiments, we observed that the use of these four loss functions and their respective weights consistently produced the best results in comparison to alternative loss combinations.

### v2

@@ Abstract
The presence of spherical distortion in equirectangular projection (ERP) images presents a persistent challenge in dense regression tasks such as surface normal estimation. Although it may appear straightforward to repurpose architectures developed for 360° depth estimation, our empirical findings indicate that such models yield suboptimal performance when applied to surface normal prediction. This is largely attributed to their architectural bias toward capturing global scene layout, which comes at the expense of the fine-grained local geometric cues that are critical for accurate surface orientation estimation. While convolutional neural networks (CNNs) have been employed to mitigate spherical distortion, their fixed receptive fields limit their ability to capture holistic scene structure. Conversely, vision transformers (ViTs) are capable of modeling long-range dependencies via global self-attention, but often fail to preserve high-frequency local detail. To address these limitations, we propose PanoNormal, a monocular surface normal estimation architecture for 360° images that integrates the complementary strengths of CNNs and ViTs. In particular, we design a multi-level global self-attention mechanism that explicitly accounts for the spherical feature distribution, enabling our model to recover both global contextual structure and local geometric details. Experimental results demonstrate that our method not only achieves state-of-the-art performance on several benchmark 360° datasets, but also significantly outperforms adapted depth estimation models on the task of surface normal prediction. The code and model are available at https://github.com/huangkun101230/PanoNormal.
keywords:
360° image; Surface normal estimation; Monocular perception.
@@ 1 Introduction
Estimating surface normal is a crucial task in computer vision as it provides essential geometric information about the structure and orientation of surfaces within a scene. By offering valuable insights into the underlying 3D geometry of objects, accurate surface normal estimation contributes significantly to advancing the capabilities of a wide range of applications, including object recognition, autonomous driving, and robotics. Despite significant progress in normal estimation for conventional perspective images [long2024adaptive, bae2021estimating, liao2019spherical], the task of 360° surface normal estimation remains less explored. When directly applying the methods proposed for perspective images to 360° images, the results are unsatisfactory (see Fig. 5). This is because the spatial warping inherent in 360° images is not taken into account in these techniques and so we instead, in this paper, develop specialized techniques tailored to their spherical representation.
To address the distortions inherent in panoramic imagery, prior work in related tasks such as 360° depth estimation has proposed projection–fusion architectures (e.g., UniFuse [jiang2021unifuse]) that fuse complementary information from equirectangular projection images and cube-map patches to produce more accurate predictions in the spherical domain. While effective for depth, these projection-fusion strategies must reconcile a domain gap between different projections and introduce significant cross-projection fusion overhead. Alternative distortion-aware approaches embed spherical geometry directly into CNNs or adopt specialized loss formulations to compensate for spatial warping [coors2018spherenet, liao2019spherical]. More recently, self-attention mechanisms have been applied to recover holistic scene structure in the 360° domain [li2022omnifusion, shen2022panoformer], with transformers demonstrating clear advantages in modelling long-range dependencies. However, despite these advances in depth estimation and global modelling, such depth-oriented and transformer-based solutions exhibit two complementary shortcomings when transferred to surface normal estimation: (1) projection-fusion and distortion-aware CNN solutions tend to prioritise correcting large-scale geometric distortions and therefore do not always preserve the high-frequency, local geometric cues required for accurate surface orientation; and (2) conventional transformer pipelines, although powerful at global reasoning, commonly lack inductive mechanisms to retain fine-grained local detail and do not explicitly exploit the spherical distribution of features. These observations — corroborated by our empirical comparisons — motivate a hybrid design that integrates CNN-based local feature extraction with multi-level, spherical-aware self-attention to jointly recover fine-scale geometry and global context for 360° surface normal estimation.
Figure 1: Our PanoNormal method produces more accurate normal estimation predictions compared to the current state-of-the-art method, particularly in the areas highlighted by the red rectangle. For better visualization, we provide a 3D point cloud generated from the ground truth depth.
@@
In this paper, we present PanoNormal, a framework for monocular surface normal estimation in 360° indoor environments. PanoNormal strategically integrates the strengths of CNNs for extracting low-level features and enhancing locality, along with the advantages of transformers in capturing and associating long-range dependencies. We introduce a deep learning module aimed at extracting low-level features from raw 360° images, departing from the conventional approach of direct tokenization inputs. To enhance the capabilities of self-attention in exploring the structure of the surface normal map, we propose a multi-scale transformer decoder, enriching the final representation by integrating information across diverse scales. As demonstrated in Fig. 1, previous state-of-the-art methods yield sub-optimal results when applied to complex scenes, failing to accurately identify complete object boundaries and providing incorrect surface normal vector directions, particularly in the areas highlighted by red boxes. In contrast, our approach predicts sharper and more accurate details of local areas and finely-delineated object boundaries, significantly improving the holistic geometry understanding of the scene. We conduct extensive experiments on public datasets to evaluate our approach, and the results demonstrate that PanoNormal consistently outperforms state-of-the-art models across all examined datasets. Our contributions are summarized as follows:
@@
  • Extensive experiments conducted on widely recognized benchmarks (namely 3D60, Stanford2D3D, Matterport3D, SunCG, and Structured3D) illustrate that PanoNormal consistently surpasses state-of-the-art approaches, and contribute to further studies in this domain by providing the first comprehensive evaluation across these public benchmarks.
@@ 2 Related Work
Most existing methods [long2024adaptive, Chen_2023_CVPR, bae2021estimating, do2020surface] for surface normal estimation are designed for perspective images, which poses challenges, such as distortions, deformations, and domain shifts, that make conventional methods unsuitable for 360° images. Moreover, panoramic images introduce additional complexities to surface normal estimation, including variable illumination and intricate indoor layouts. Recent works can be broadly categorized into direct and indirect methods. Direct methods [liao2019spherical, karakottas2019360, feng2020deep] introduce specific loss functions (e.g., spherical loss, hyper-sphere loss, double-quaternion loss) to comprehend the geometric structure on the unit sphere with various CNN architectures. Coors et al. [coors2018spherenet] explicitly embed invariance against spherical distortions into CNNs, adjusting the sampling locations of convolutional filters to counter distortions and align the filters around the sphere effectively. Indirect methods [albanis2021pano3d, zhao2022monovit] draw inspiration from other vision tasks, e.g., 360° depth estimation, and adopt their architecture for surface normal estimation. Other work [li2022omnifusion, jiang2021unifuse, wang2022bifuse++, shen2022panoformer] proposes methodologies that project equirectangular images into corresponding cube map or tangent plane projection sub-images for predictions or directly learn internal correlations among these projections, aiming to capture details and holistic information concurrently with distortion awareness. However, employing simple CNNs with various spherical loss functions, adopting surface normal estimation methods from depth alone, or calculating surface normals directly from depth maps [long2024adaptive] fails to yield satisfactory results. Besides, other studies have explored complementary directions in 360° vision. Huang et al. [huang2023360] and Kou et al. [kou2024neural, kou2025omniplane] introduced image editing operations specifically designed for 360° representations. Wang et al. [wang2025target] investigated target scanpaths to enhance 360° media experiences, while Peng et al. [peng2025robust] utilized dynamic gnomonic projection for object tracking in the 360° domain. Although these works broaden the scope of 360° research, they remain distinct from surface normal estimation. In addition, to the best of our knowledge, there are no surface normal estimation methods that provide a standardized evaluation across diverse public benchmarks, as has been established in depth estimation tasks.
@@ 2.2 Vision Transformer
Vision transformers are a class of models that use self-attention mechanisms to process images as sequences of patches [dosovitskiy2020image]. ViTs [zhao2022monovit, wang2022uformer, jain2023semask, Zhu_2023_CVPR] have achieved remarkable results on conventional images captured by cameras that have a limited field of view and that have, at most, mild distortions. In contrast, 360° images offer a complete, immersive view of the surrounding scene but require adaptations and modifications of the original ViTs to handle challenges such as significant image distortions, object deformation, and domain shifts. Shen et al. [shen2022panoformer], Li et al. [li2022omnifusion], Yun et al. [yun2023egformer], Zhang et al. [zhang2025sgformer], and Ai et al. [ai2023hrdfuse, ai2024elite360d] proposed extracting features from other projection domains, such as cubemap, tangent, and icosahedron domains, to complement the ERP domain and reduce the negative effects of panoramic distortions, and introduced an attention module that considers both spatial and angular relations among tokens. Yun et al. [yun2022panoramic] presented a panoramic ViT for saliency detection in 360° videos, which leverages a multi-scale feature fusion module and a temporal attention module to capture the spatial and temporal saliency cues. Zhang et al. [Zhang_2022_CVPR] proposed a method for panoramic semantic segmentation task, which is equipped with deformable patch embedding and deformable MLP modules for handling object deformations and image distortions, and also enhances the mutual prototypical adaptation strategy for unsupervised domain adaptive panoramic segmentation. Recently, multi-task learning for 360° images has gained attention for jointly estimating depth and surface normals. Huang et al. proposed the 360MTL [huang2025multi], which employs a spherical-aware Vision Transformer [shen2022panoformer] to handle distortion and achieved strong performance on both tasks. These works demonstrate the potential and effectiveness of ViTs for 360° vision tasks and motivated us to investigate suitable ViTs tailored for the surface normal dense prediction task, aiming to seamlessly integrate the strengths of the previously mentioned methods. Our objective is to yield robust evaluation results across benchmarks, contributing to further studies in this domain.
@@ 3 Our Method
To address the challenge of estimating surface normals using a single panoramic image for indoor scenes, we propose a novel architecture, PanoNormal. PanoNormal leverages the capabilities of convolutional layers for extracting meaningful embeddings from panoramic images. To address the challenge of spherical distortion in 360° images while maintaining global dependencies, we crafted a U-shaped distortion-aware transformer architecture, drawing inspiration from the approach proposed by Shen et al. [shen2022panoformer]. More specifically, we adopt a multi-level structure in our transformer decoder, capturing spatial relationships, processing fine-grained details, and integrating high-level contextual information across different levels. Fig. 2 illustrates the architecture of PanoNormal.
@@ 3.1.1 Feature Embedding
In a conventional ViT-based network, feature embedding undergoes processing through only a single convolutional layer before ViT blocks. It has been demonstrated that this works well for some 360° understanding tasks, such as depth estimation [shen2022panoformer]. Nonetheless, although both depth estimation and surface normal tasks demand high-quality contextual information, it is not effective to apply a single convolutional layer to tackle both depth and surface normal estimation tasks. This is primarily due to the intricate nature of these tasks and their feature requirements. Depth estimation benefits from a broader, more global understanding of the scene, which can be achieved with a single convolutional layer. In contrast, surface normal estimation requires capturing detailed local structures and fine patterns, which is better accomplished with multiple convolutional layers. To address this issue, we introduce a series of convolutional layers aimed at extracting more effective features. This approach enhances our model’s ability to gain a more comprehensive understanding of the entire scene, improving performance in surface normal estimation. The feature embedding block consists of three $3\times 3$ convolutional layers, each followed by a batch normalization function and a rectifid linear unit (ReLU) activation function.
@@ 3.1.2 Distortion-aware Transformer Encoder
The inherent non-uniform spatial warping of visual features, caused by spherical distortion in the 360° image, is built upon the PanoFormer transformer mechanism [shen2022panoformer]. To address issues such as distortions, misalignment, and inaccuracies arising from direct pixel sampling on the ERP image, a tangent projection division strategy is employed. It samples related surrounding pixels of the central pixel (tangent point) on each dense divided plane in the tangent domain, obtaining a series of tokens in patches with decent position representation that can be further transformed to the ERP domain. Subsequently, a conventional multi-head vision transformer block is applied, with the replacement of the feed-forward network to the locally-enhanced feed-forward network [yuan2021incorporating, wang2022uformer] for enhancing local feature interaction. The introduced encoder not only computes attention scores between central and correlated tokens but also integrates a trainable token flow. This token flow serves as a bias to adjust the spatial distribution of tangent tokens, supplying the network with additional positional information for learning global dependencies. An example is shown in Fig. 2(a) and the representation of the self-attention is as follows:
@@
where the sampling strategy is adopted on the feature representations $f$, and is denoted as $\hat{s}$. The parameters $m,q,k$ represent the self-attention head, each token, and its neighboring tokens in a tangent patch, respectively. $W_{m}$ and $W^{\prime}_{m}$ denote the learnable weights of each head, $A_{mqk}$ is the attention weights for each token, and $\Delta s_{mqk}$ signifies the learned bias of each token’s query.
@@ 3.1.3 Multi-level Transformer Decoder
Our proposed ViT decoder, designed with spherical distortion awareness and a hierarchical structure to predict normal maps of different scales, is illustrated in Fig. 2(b). The proposed multi-level decoder comprises four independent blocks. Each block processes the concatenation of the upsampled encoded representations $\hat{f}_{i}$, featuring twice the spatial resolution size and half the number of channels from the preceding level, and the directly propagated features $\hat{s}_{i}$ from the encoders through the skip-links. The surface normal vectors in various scales are generated as outputs using $3\times 3$ convolutions, followed by a hyperbolic tangent (tanh) activation function to constrain them to the range [1,1][-1,1]. This process can be described as:
@@ 3.2 Loss Function
The proposed network generates surface normal maps at multiple scales to capture the scene’s detailed global geometric structure. During training, the predicted maps are upsampled to match the input size using bilinear interpolation. The training process incorporates MSE, quaternion, perceptual, and smooth loss functions, which are detailed as follows:
MSE Loss, $L_{m}$ is the mean squared error between the ground truth and the predicted normals of each pixel, defined as:
@@
Quaternion Loss, $L_{q}$ [karakottas2019360] measures the angular difference between predicted and ground truth normal maps on a pixel-wise basis:
@@
where $M$ indicates the number of pixels of the input, and $j$ indexes the current pixel. Perceptual Loss, $L_{p}$ is employed on the finest scales to enhance the generation of finer details:
$L_{p}=l_{\textsubscript{feat}}^{\phi,k}(\hat{\textbf{N}},\textbf{N})=\sum_{j=1}^{M}\frac{1}{C_{k}M}\|\phi_{k}(\hat{\textbf{N}}_{j})-\phi_{k}(\textbf{N}_{j})\|_{2}^{2}$ (5)
where $\phi$ is the VGG-16 network [simonyan2014very] that pretrained on the ImageNet dataset [russakovsky2015imagenet], and $C_{k}$ indicates $C$ dimensional features for the $k$-th layer of the network $\phi$.
Smooth Loss, $L_{s}$ quantifies the gradient, $G$, in the $x$ and $y$ directions in the ground truth and the predicted surface normal map at all scales:
@@
By default, we set $\lambda_{m}=1.0$, $\lambda_{q}=10.0$, $\lambda_{p}=0.05$, and $\lambda_{s}=0.5$ as the weights for different terms. Our experiments show that this specific combination consistently yields the best results compared to other loss configurations. The effectiveness of each loss function is further validated in our ablation study in Sec. 4.4.2